OpenStudy (anonymous):

integral of (x^2+1)/(2x)dx from 1 to 2

OpenStudy (ikram002p):

Hiint :- \(\dfrac{x^2+1}{2x}=\dfrac{ 1}{2x}+\dfrac{x^2 }{2x}\)

OpenStudy (unklerhaukus):

\[\int_1^2\frac{x^2+1}{2x}\,\mathrm dx\\ =\frac12\int_1^2\frac{x^2+1}{x}\,\mathrm dx\\ =\frac12\left(\int_1^2\frac{x^2}{x}\,\mathrm dx+\int_1^2\frac{1}{x}\,\mathrm dx\right)\\ =\frac12\left(\int_1^2x\,\mathrm dx+\int_1^2\frac{1}{x}\,\mathrm dx\right)\\=\]